AAVSOnet news, 2015-02-20

Fri, 02/20/2015 - 12:51

SRO is currently down.  There is a dome issue, and John Gross is/has been out of town.  He will look into it as soon as he returns, which should be shortly.  It hasn't made any impact, however, because the weather has been poor since the failure.  It does not look to improve for several more days.

The Celectron CGEM mounts that we use for the BSM systems are great except for one major remote-observing flaw.  If you have a power failure, the mount reverts back to hand-paddle control, where you must do a manual alignment before control can be returned to the computer.  This means a human has to physically be at the observatory to fix things.  BSM_NM ran into this a couple of days ago, where a daytime power failure occurred.  The technician is due to arrive at the observatory on Saturday to restore operation.  Again, the southwest has poor weather so no loss of observing has occurred.

TMO61 remains down until Jon Holtzman returns from his workshop.  I expect that operations will resume in about a week.

OC61 remains down until a replacement motor controller arrives at Mt. John and is installed.  That controller is now on its way to New Zealand, so with luck, OC61 will be back on line next week.  During bright time, we will be running some tests of the new spectrograph, and hope to make it available to the community in March.

BSM HQ is down because of snow on the HQ roof.  I will look at it during my next visit to HQ and see if we can safely use the system.

Lots of current "down" systems!  So what IS currently working?

BSM Hamren and BSM South are fully operational.  Remember that BSM South has a diffraction grating (donated by Frank Schorr), and we have the system programmed to be taking grating images of both of the new novae.

BSM Berry is now operational.  We are in the process of setting up the file transfer to HQ and modifying the pipeline to include this system, but it looks great.  Berry has a 10-position filter wheel, including BVRI, g'r'i'z', and two diffraction gratings rotated 90 degrees from one another.

Coker30 is working great, now that the cable issue has been resolved.  Bill Goff says that he will be putting the diffraction grating donated by Frank Schorr into the filter wheel soon.

As you can see, most of the telescopes that are off-line right now should be operating within a week or so.